OpenStudy (anonymous):

simplify the radical -4sqrt 6x^5y^6 * 7 sqrt 8x^3y show work

OpenStudy (anonymous):

\[-4\sqrt{6x^5y^6}*7\sqrt{8x^3y}\]

OpenStudy (anonymous):

Is that the question ^^^ ?

OpenStudy (anonymous):

yes that is the question

OpenStudy (anonymous):

when multiplying radicals, you can combine them like this: \[\sqrt{a}*\sqrt{b}=\sqrt{a*b}\]

OpenStudy (anonymous):

ok and how do i do that if they are different

OpenStudy (anonymous):

so, we can rewrite as below \[-4*7*\sqrt{6*8*x ^{5+3}*y ^{6+1}}\]

OpenStudy (anonymous):

is the above clear ? what do you mean "if they are different" ?

OpenStudy (anonymous):

ok i see what you did now what do i do next

OpenStudy (anonymous):

next: rewrite as below: \[-28*\sqrt{3*4^2*x^8*y^7}\]

OpenStudy (anonymous):

so it should read \[-28\sqrt{48x^8y^7}\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

nevermind what i got i just multiplied everything

OpenStudy (anonymous):

ok what do i do after i get what you have

OpenStudy (anonymous):

next \[-28*4*\sqrt{3}*\sqrt{x^8}*\sqrt{y}*\sqrt{y^6}\]

OpenStudy (anonymous):

what do i do after that

OpenStudy (anonymous):

\[-112*\sqrt{3}*x^4*y^3*\sqrt{y}\]

OpenStudy (anonymous):

that's about as far as you can go here.

OpenStudy (anonymous):

you could also write it like this: \[-112*\sqrt{3}*x^4*y ^{\frac{7}{2}}\]

OpenStudy (anonymous):

Is it all clear ? If not let me know where it is not :)

OpenStudy (anonymous):

not sure where you got-112 from

OpenStudy (anonymous):

from -28*4

OpenStudy (anonymous):

oh ok and what about the other things

OpenStudy (anonymous):

Remember we had this number inside the radical ? \[\sqrt{48}=\sqrt{3*4^2}=\sqrt{3}*\sqrt{4^2}=\sqrt{3}*(4^{2})^{\frac{1}{2}}=\sqrt{3}*4\] the -112 came from: -7*4 (which were always outside the radicals) * 4 (that's the 4 from above, taken out of the radical)

OpenStudy (anonymous):

ok what about \[\sqrt{3}*x^4*y^3*\sqrt{y}\] what do i do with these ad how did you get them

OpenStudy (anonymous):

you can't do much more with these. The way I got them, is in the history of posts I made above. We can review them one by one, to make sure they are all clear. If you want to remove radicals completely from the expression you can convert like this: \[\sqrt{x}=x ^{\frac{1}{2}}\] Do yo want to do that ?

OpenStudy (anonymous):

can we go through each separately

OpenStudy (anonymous):

sure. The first thing was to combine the two radicals into one long one (see my 3rd and 4th post above).

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

If you look at the expression in the 4th post, you will see that x and y have powers that are sums of two numbers. Do you see why that is ?

OpenStudy (anonymous):

i dont really understand it

OpenStudy (anonymous):

How much is: x^2*x^3=?

OpenStudy (anonymous):

x^5

OpenStudy (anonymous):

right, x^2*x^3=x^(2+3)=x^5 And how much is: x^5*x^3=?

OpenStudy (anonymous):

x^8

OpenStudy (anonymous):

how much is: \[\sqrt{x^5}*\sqrt{x^3}\]

OpenStudy (anonymous):

hint: combine into one radical.

OpenStudy (anonymous):

\[\sqrt{x^3}*x^5\]

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

It should be like this: \[\sqrt{x^5}*\sqrt{x^3}=\sqrt{x^5*x^3}=\sqrt{x ^{5+3}}=\sqrt{x^8}=(x^8)^{0.5}=x^{8*0.5}=x^4\]

OpenStudy (anonymous):

I broke it into lots of steps on purpose ... go over it and let me know if it is all clear.

OpenStudy (anonymous):

there are 7 steps above. let me know if any of them is unclear.

OpenStudy (anonymous):

what is after \[(x^8)^{0.5}=x^8\] i cant see it

OpenStudy (anonymous):

The last part is: \[(x^8)^{0.5}=x^{8*0.5}=x^4\]

OpenStudy (anonymous):

can i email you a link and you actually show me how we are doing this because i think i have confused myself

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

post the link here

OpenStudy (anonymous):

ok here it is http://www.twiddla.com/605550

OpenStudy (anonymous):

no unwanted people please